# MuSeD: A Multimodal Spanish Dataset for Sexism Detection in Social Media Videos

### v1

@@ 4.3 Results
Claude-3.7 Sonnet 82.00 82.50
Gemini-2.0-Flash (Video) 67.17 68.92
@@
The results in Table 3 suggest that models can detect sexism to some extent, with accuracy varying across different model sizes. Larger models generally outperform smaller ones. The best-performing model in this setting is GPT-4o, achieving 80.50% accuracy, followed by Claude 3.7 Sonnet and Qwen2.5-32B-Instruct (76.50%). The gap between models suggests that sexism detection remains a challenging task, especially for models with fewer parameters where alternative approaches such as few-shot prompting or task-specific training may be necessary to enhance sexism detection without requiring larger models.
@@
Comparing results from text-only inputs with text and visual content (Table 3), we observe a performance increase in some models. GPT-4o achieves 83.50% accuracy, while Claude 3.7 Sonnet achieves 82.50%. However, not all models benefit equally from visual information. For instance, Gemini 2 Flash performs worse with raw video input (68.92%) compared to preprocessed frames (72.50%), suggesting that data-specific preprocessing techniques like our frame extraction method (see Section 4.2) may enhance the quality of visual cues. This variation highlights that while visual information can improve performance, its effectiveness depends on how well a model integrates and uses multimodal data.
Can text-only models fully capture sexism?
Text-only models perform slightly worse when evaluated against the multimodal label, indicating that textual cues alone do not fully capture sexism when visual context is considered by annotators. For example, the performance of GPT-4o on the multimodal label, 79.5%, is slightly below its accuracy of 80.5% on the text label. On the other hand, multimodal models perform well against the text label, with GPT-4o and Claude-3.7 Sonnet outperforming text-only models (by 1.5 and 5.5 points respectively), indicating that they effectively integrate textual signals. These results confirm the idea that multimodal models successfully incorporate text but also rely on visual information, whereas text-only models may overlook key visual signals that influence human judgments.
@@
We also analyze the 13 cases where videos contain primarly background music, meaning the main textual information comes from OCR. GPT-4o achieves 84.6% accuracy on the text label when using text input alone but drops to 54.8% on the multimodal label when incorporating both text and images. Similarly, Claude 3.7 Sonnet achieves 76.9% accuracy with text-only input and 61.5% with multimodal input. These results highlight the critical role of textual information, even when derived solely from OCR, in detecting sexism.
How do the models perform in detecting sexism in videos from BitChute?
In the same way as for the annotation task, we expect the models to correctly classify a consistent set of videos collected from BitChute as sexist (see Section 3.4). Using GPT-4o with the text label, we find that the model correctly classifies 81.82% of the cases as sexist. When GPT-4o is used with both text and video, it correctly classifies 75.76% of the cases as sexist. We find the opposite situation using Claude 3.7 Sonnet. With the text label, the model classifies 69.70% of the cases correctly, and when it is given access to text and video, the accuracy score increases to 93.94%. This aligns with Table 3, where Claude 3.7 Sonnet’s performance improves when provided with visual information.
@@ Acknowledgments
This work was supported by the ‘FairTransNLP-LANGUAGE: Analyzing toxicity and stereotypes in language for unbiased, fair, and transparent systems’ project (MCIN/AEI/10.13039/501100011033/FEDER,UE) and by a research grant (VIL53122) from VILLUM FONDEN. We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina, hosted by IT4Innovations, Czech Republic.
@@
The primary intent of this work is to advance research in the automatic detection of harmful content on social media platforms. Our goal is to contribute to the development of AI-driven moderation systems that help create a fair and safe online environment while also supporting human moderators, who are at risk of mental health issues due to prolonged exposure to harmful content. To ensure ethical and unbiased data annotation, we carefully recruited annotators with diverse gender and age backgrounds, minimizing annotation bias in the labeling process. All annotators provided informed consent to engage with potentially offensive material. Additionally, we conducted weekly discussions where annotators could openly share their thoughts and emotional responses, with a moderator ensuring a supportive and respectful environment. The annotators were paid €10 per hour (50 hours in total). To uphold ethical standards, we manually excluded videos containing explicit violent content and only collected publicly available videos shared by their creators.
@@ Appendix A Annotation guidelines
The focus of the annotation is to classify content as either sexist or non-sexist. If the content is deemed sexist, the specific segment containing sexism must be selected. Content is considered sexist if it includes:
@@
Content that discriminates against individuals based on their gender identity, including criticism of the ability to choose a gender different from the sex assigned at birth A, as seen in the case of transgender individuals.
@@ Appendix B Annotator profiles
Table 4 shows the annotator profiles, including gender, age, native language, study, and occupation.
@@
Study/Occupation Consultant (1), Student (2),
Teacher (2), Adjunct prof. (1)
@@ Appendix D Corpus statistics
Table 5 shows dataset statistics, Figure 3 shows the distribution of video lengths and Figure 4 the frequency of the top 10 hashtags used.
@@ Appendix F Evaluation prompts
Figure 5: Prompt used to classify text transcripts from videos using text-only models. The positions are indicated by placeholders [$\ldots$], where the text to be classified is inserted.
@@
Figure 6: Prompt used to classify videos using Multimodal LLMs. The positions are indicated by placeholders [$\ldots$], where the text to be classified is inserted. The images are also given to the models.
@@
Figure 7: Prompt used to classify videos using Gemini-2.0=Flash (Video). The positions are indicated by placeholders [$\ldots$], where the text to be classified is inserted. The video is also given to the model.
@@ F.1 Alternative prompts
In the case of Qwen2.5-32B-Instruct, we find the best results using an analogous text prompt translated to english. The results in accuracy for the text label increase to 79.13% and to 78.36% in the case of the multimodal label. This great improvement is likely due to the model train set containing more english language tokens.
@@
Figure 8: English prompt used to classify text transcripts from videos using text-only models. The positions are indicated by placeholders [$\ldots$], where the text to be classified is inserted.

### v2

@@ 4.3 Results
Claude-3.7 Sonnet 82.50 82.00
Gemini-2.0-Flash (Video) 68.92 67.17
@@
The results in Table 3 suggest that models can detect sexism to some extent, with accuracy varying across different model sizes. [^20] Table 9 reports LLMs performance broken down by dataset source, including accuracy for both TikTok and BitChute, as well as the difference between them. Larger models generally outperform smaller ones. The best-performing model in this setting (text label) is GPT-4o, achieving 80.50% accuracy, followed by Claude 3.7 Sonnet and Qwen2.5-32B-Instruct (76.50%). The gap between models suggests that sexism detection remains a challenging task, especially for models with fewer parameters where alternative approaches such as few-shot prompting or task-specific training may be necessary to enhance sexism detection without requiring larger models.
Label
Prompt
Text
Basic
Text
Definition
Multimodal
Basic
Multimodal
Definition
Random 50.00 50.00 50.00 50.00
Llama-3-70B-Instruct 71.50 83.25 69.42 82.25
Gemini-2.0-Flash 68.25 75.25 67.25 72.75
GPT-4o 80.50 82.50 79.50 81.00
Claude-3.7 Sonnet 76.50 84.25 75.00 83.75
Vision & Language
Gemini-2.0-Flash (Video) 68.92 77.25 67.17 75.75
Gemini-2.0-Flash 72.50 74.00 72.50 74.00
GPT-4o 83.50 86.50 82.00 85.00
Claude-3.7 Sonnet 82.50 87.75 82.00 86.75
Table 4: Accuracy comparison across models using basic (yes/no) and definition prompts, evaluated against text and multimodal labels. Best results in each group are in bold.
@@
Comparing results from text-only inputs with text and visual content (Table 3 Text Label), we observe a performance increase in some models. GPT-4o achieves 83.50% accuracy, while Claude 3.7 Sonnet achieves 82.50%. However, not all models benefit equally from visual information. For instance, Gemini 2 Flash performs worse with raw video input (68.92%) compared to preprocessed frames (72.50%), suggesting that data-specific preprocessing techniques like our frame extraction method (see Section 4.2) may enhance the quality of visual cues. This variation highlights that while visual information can improve performance, its effectiveness depends on how well a model integrates and uses multimodal data.
How do LLMs perform against multimodal annotations?
Models perform slightly worse when evaluated against the multimodal label compared to the text label, suggesting that incorporating visual context introduces additional complexity. This pattern holds across both text-only and multimodal models. For instance, GPT-4o achieves 83.5% accuracy on the text label but drops to 82% on the multimodal label. Similarly, Claude-3.7 Sonnet sees a decrease from 82.5% to 82%. These differences indicate that visual signals may alter human judgments in ways that are harder for current models—particularly those trained primarily on text—to fully capture. Nonetheless, multimodal models still perform competitively across both settings, and their strong performance on the text label (e.g., GPT-4o outperforming all text-only models by at least 3 percentage points) suggests they effectively integrate textual information. Overall, the results highlight that the multimodal label introduces a more challenging benchmark and underscore the importance of visual context in shaping perceptions of sexism.
Do LLMs benefit from a more detailed definition of sexism?
As shown in Table 4, incorporating the definition consistently improves model performance across both text-only and multimodal settings. For the text label, models such as Llama 3 and Claude 3 show notable improvements, with accuracy gains from 71.50% to 83.25% and from 76.50% to 84.25%, respectively. Similarly, for the multimodal label, GPT-4o and Claude 3 achieve accuracy improvements from 82.0% to 85% and 86.75% respectively. These results show that providing a more detailed definition enhances not only human annotation quality but also the model’s ability to accurately identify sexist content, underscoring the importance of precise task formulation.
@@
We also analyze the 13 cases where videos contain primarily background music, meaning the main textual information comes from OCR. GPT-4o achieves 84.6% accuracy on the text label when using text input alone, but drops to 54.8% on the multimodal label when incorporating both text and images. Similarly, Claude 3.7 Sonnet achieves 76.9% accuracy with text-only input and 61.5% with multimodal input. These results highlight the critical role of textual information, even when derived solely from OCR, in detecting sexism.
How do LLMs perform in detecting sexism in videos from BitChute?
In the same way as for the annotation task, we expect the models to correctly classify a consistent set of videos collected from BitChute as sexist (see Section 3.4). GPT-4o reaches 88% accuracy on the text label when using only textual input, but its performance decreases to 82% on the multimodal label when both text and images are used. In contrast, Claude 3.7 Sonnet shows the opposite trend: it achieves 70% accuracy with text-only input, which rises to 94% when incorporating multimodal input. This aligns with Table 3, where Claude 3.7 Sonnet’s performance improves substantially when provided with visual information.
@@ Acknowledgments
This work was supported by the ‘FairTransNLP-LANGUAGE: Analyzing toxicity and stereotypes in language for unbiased, fair, and transparent systems’ project (MCIN/AEI/10.13039/501100011033/FEDER,UE), and funded by the CLiC research group (2021SGR00313) and a research grant (VIL53122) from VILLUM FONDEN. We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina, hosted by IT4Innovations, Czech Republic.
@@
The primary intent of this work is to advance research in the automatic detection of harmful content on social media platforms. Our goal is to contribute to the development of AI-driven moderation systems that help create a fair and safe online environment while also supporting human moderators, who are at risk of mental health issues due to prolonged exposure to harmful content. To ensure ethical and unbiased data annotation, we carefully recruited annotators with diverse gender and age backgrounds, minimizing annotation bias in the labeling process. All annotators provided informed consent to engage with potentially offensive material. Additionally, we conducted weekly discussions where annotators could openly share their thoughts and emotional responses, with a moderator ensuring a supportive and respectful environment. The annotators were paid €10 per hour (50 hours in total). To uphold ethical standards, we manually excluded videos containing explicit violent content and only collected publicly available videos shared by their creators. None of the authors directly participated in the annotation process.
@@ Appendix A Annotation guidelines
The focus of the annotation is to classify content as either sexist or non-sexist. Content is considered sexist if it includes:
@@
Content that discriminates against individuals based on their sexual orientation and gender identity, including criticism of the ability to choose a gender different from the sex assigned at birth A, as seen in the case of transgender individuals.
@@ Appendix B Annotator profiles
Table 5 shows the annotator profiles. The annotators were recruited through academic channels and all have a background in linguistics. Two are current undergraduate students pursuing a degree in linguistics, one is an adjunct professor specializing in linguistics, and the remaining annotators were selected based on prior academic experience in data annotation and corpus creation. All annotators received comprehensive training before beginning the annotation process to ensure consistency and reliability.
@@
Study/Occupation Linguistics consultant (1), Linguistics students (2),
Teacher (2), Adjunct Professor in Linguistics (1)
@@ Appendix D Corpus statistics
Table 6 presents the dataset statistics, while Table 7 shows the distribution of sexist and non-sexist videos across TikTok and BitChute. Figure 4 shows the distribution of the video lengths, and Figure 4 shows the frequency of the top 10 hashtags used.
@@
Platform Number of Videos Sexist Cases Non-Sexist Cases
TikTok 367 164 203
BitChute 33 30 3
Table 7: Distribution of sexist and non-sexist videos on TikTok and BitChute.
@@ Results
Table 9 reports LLMs and Multimodal LLMs accuracy by dataset source. Smaller models tend to exhibit a wider performance gap between the two data sources, while larger models show more consistent performance.
Model BitChute Accuracy TikTok Accuracy Difference
LLMs (Text Label)
Llama-3-8B-Instruct 0.12 0.55 0.43
Qwen2.5-3B-Instruct 0.42 0.63 0.21
Salamandra-7b-instruct 0.70 0.58 -0.12
Llama-3-70B-Instruct 0.42 0.74 0.32
Gemini-2.0-Flash 0.64 0.69 0.05
Qwen2.5-32B-Instruct 0.70 0.77 0.07
GPT-4o 0.88 0.80 -0.08
Claude 3.7 Sonnet 0.70 0.77 0.07
Multimodal LLMs (Multimodal Label)
GPT-4o 0.82 0.82 0.00
Gemini-2.0-Flash 0.79 0.72 -0.07
Claude 3.7 Sonnet 0.94 0.81 -0.13
Gemini-2.0-Flash (Video) 0.70 0.67 -0.03
Table 9: Accuracy of LLMs and Multimodal LLMs for sexism detection across BitChute and TikTok.
@@ Appendix F Evaluation prompts
Figure 5: Prompt used to classify text transcripts from videos using text-only models. The placeholder $\ldots$] indicates where the text to be classified is inserted.
@@
Figure 6: Prompt used to classify videos using Multimodal LLMs. The placeholder $\ldots$] indicates where the text to be classified is inserted. The images are also given to the models.
@@
Figure 7: Prompt used to classify videos using Gemini-2.0=Flash (Video). The placeholder $\ldots$] indicates where the text to be classified is inserted. The video is also given to the model.
@@ F.1 Alternative prompts
In the case of Qwen2.5-32B-Instruct, we find the best results using an analogous text prompt translated to English. The text label accuracy increase to 79.13% and to 78.36% in the case of the multimodal label. This great improvement is likely due to the model train set containing more English language tokens.
@@
Figure 8: English prompt used to classify text transcripts from videos using text-only models. The positions are indicated by placeholders $\ldots$], where the text to be classified is inserted.
F.2 Extended prompts
System prompt
User prompt
### Instrucción:
Determina si el texto es sexista. A continuación se presentan criterios que definen el sexismo:
Estereotipos: (1) Formulación de propiedades descriptivas que supuestamente distinguen a hombres y mujeres basadas en estereotipos de género. (2) Formulación de propiedades prescriptivas que hombres y mujeres deben cumplir para encajar en los roles de género establecidos por la sociedad. Desigualdad: (3) Contenido que niega la existencia de desigualdades (pasadas o presentes) entre hombres y mujeres. (4) Contenido que se opone al feminismo, argumentando que este movimiento margina a los hombres. Discriminación: (5) Contenido que discrimina a personas por su orientación sexual (por ejemplo, personas homosexuales, lesbianas, intersexuales, bisexuales, asexuales y pansexuales), por su identidad de género, incluyendo críticas hacia quienes eligen un género distinto al asignado al nacer (por ejemplo, personas transgénero). También contenido que discrimina a personas LGBTQ+. Cosificación: (6) Contenido que presenta a las mujeres como objetos, evalúa su apariencia física o las critica por no ajustarse a estándares normativos de belleza. ### Text:
$\ldots$]
### ¿Es el texto sexista?:
Figure 9: Extended prompt used to classify text transcripts from videos using text-only models. The placeholder $\ldots$] indicates where the text to be classified is inserted.